# Multiscale Training of Convolutional Neural Networks

### v1

@@ 5 Experimental Results
In this section, we evaluate the empirical performance of our training strategies: Multiscale-SGD (Algorithm 1) and Full-Multiscale-SGD (Algorithm 2), compared to the standard CNN training with SGD. We demonstrate these strategies on architectures like ResNet (He et al., 2016) and UNet (Ronneberger et al., 2015), and their MFC versions, where we replace standard convolutions with our mesh-free convolutions (Section 3.1). We denote these modified architectures as MFC-X, where X is the original network architecture (e.g., MFC-ResNet and MFC-UNet). Additional details on experimental settings, hyperparameters, architectures, and datasets are provided in Appendix E.
@@
Metrics. To address our questions, we focus on performance metrics (e.g., MSE or SSIM) and the computational effort for each method. As a baseline, we use a standard ResNet or UNet trained on a single input resolution. We define a work unit (#WU ) as one application of the model on a single image at the finest mesh (i.e., original resolution). In multiscale training, this unit decreases by a factor of 4 with each downsampling. We then compare #WU  across training procedures.
Blur Estimation. We begin by solving a quadratic problem to validate the theoretical concepts discussed in Sections 2.2 and 3.2. Specifically, we estimate a blurring kernel directly from data (see Nagy & Hansen (2006) for the definition). We assume both the image ${\bf u}$ and its blurred version ${\bf y}$ are available on a $256^{2}$ fine mesh. The relationship is linear, given by ${\bf y}={\bf G}\star{\bf u}+\epsilon$, where ${\bf G}$ is the blurring kernel. Given ${\bf u}$ and ${\bf y}$, we solve the optimization problem $\min_{{\bf G}}{\frac{1}{2}}{\mathbb{E}}_{{\bf u},{\bf y}}\|{\bf y}-{\bf G}\star{\bf u}\|^{2}.$
This is a quadratic stochastic optimization problem for the kernel ${\bf G}$. We simulate ${\bf u}$ as a Random Markov Field (Tenorio et al., 2011) using a random $7\times 7$ blurring kernel for ${\bf G}$. The problem is solved with standard CNNs and MFCs, comparing SGD, Multiscale-SGD, and Full-Multiscale-SGD strategies. As seen in Figure 4, for standard convolutions, the loss increases with mesh refinement, making multiscale approaches ineffective. However, our MFCs maintain stable loss across meshes, making the multiscale process efficient, with the final 50 iterations on the finest mesh. We measure computational work by running 200 iterations on the fine mesh with SGD, Multiscale-SGD, and Full-Multiscale-SGD, and early-stopping when the test loss plateaus. Table 3 shows that using Multiscale-SGD yields significant savings, and Full-Multiscale-SGD offers even further savings.
(a) Standard Convolution
(b) Mesh-Free Convolution
(c) Training Loss
@@
Image Denoising. Here, one assumes data of the form ${\bf u}^{h}=t{\bf y}^{h}+(1-t){\bf z}$ where ${\bf y}^{h}$ is some image on a fine mesh $h$ and ${\bf z}\sim N(0,{\bf I})$ is the noise. The noise level $t\in[0,1]$ is chosen randomly. The goal is to recover ${\bf u}^{h}$ from ${\bf y}^{h}$. The loss to be minimized is $loss({\bm{\theta}})={\frac{1}{2}}{\mathbb{E}}_{{\bf y}^{h},{\bf u}^{h},t}\|f({\bf u}^{h},t,{\bm{\theta}})-{\bf y}^{h}\|^{2}.$ We use the STL10 dataset, and the training complexity of SGD with Multiscale-SGD from Algorithm 1, and Full-Multiscale-SGD from Algorithm 2 on several architectures. The results are presented in Table 4, with additional results on the CelebA dataset in Table 6. We notice that both multiscale training strategies achieved similar performance to SGD for all networks (see Table 4 and Table 6) with a considerably lower number of #WU . The calculations for #WU for the three training strategies have been presented in Appendix C. We provide visualizations of the results in Appendix G, and additional details on the experiment in Appendix E.
Training Strategy #WU ($\downarrow$) MSE ($\downarrow$)
UNet ResNet MFC-UNet MFC-ResNet
SGD (Single Scale) 480,000 0.1918 0.1629 0.1800 0.1862
@@
Table 4: Comparison of training strategies on the denoising task on the STL10 dataset.
Image Super-Resolution. Here we aim to predict a high-resolution image ${\bf u}^{h}$ from a low-resolution image ${\bf y}^{l}$, which is typically a downsampled version of ${\bf u}^{h}$. The downsampling process is modeled as ${\bf y}^{l}=D{\bf u}^{h}+{\bf z}$, where $D(\cdot)$ is a downsampling operator (e.g., bicubic), and ${\bf z}$ is noise. The goal is to invert this process and reconstruct ${\bf u}^{h}$ using a neural network $f({\bf y}^{l},{\bm{\theta}})$. The loss function is $\text{loss}({\bm{\theta}})=\frac{1}{2}\mathbb{E}_{{\bf y}^{l},{\bf u}^{h}}\left\|f({\bf y}^{l},{\bm{\theta}})-{\bf u}^{h}\right\|^{2}.$ The model $f({\bf y}^{l},{\bm{\theta}})$ predicts the high-resolution image ${\bf u}^{h}$ from the low-resolution input ${\bf y}^{l}$.In Table 5, our Full-Multiscale-SGD significantly accelerates training while maintaining image quality, measured by SSIM. Additional details and visualizations are provided in Appendix E and Appendix G.
Training Strategy #WU ($\downarrow$) SSIM ($\uparrow$)
@@
SGD (Single Scale) 16,000 0.84 0.82 0.84 0.83
Multiscale-SGD (Ours) 2,500 0.82 0.81 0.83 0.83
Full-Multiscale-SGD (Ours) 340 0.81 0.81 0.82 0.83
Table 5: Comparison of SSIM for ESPCN (Shi et al., 2016), ResNet, MFC-ESPCN and MFC-ResNet on the Urban100 dataset for the Super-Resolution using different training strategies.
6 Conclusions
In this paper, we introduced a novel approach to multiscale training for deep neural networks, addressing the limitations of standard mesh-based convolutions. We showed that theoretically, standard convolutions may not converge as $h\rightarrow 0$ for noisy inputs, hindering performance. To solve this, we proposed mesh-free convolutions inspired by parabolic PDEs with trainable parameters. Our experiments demonstrated that mesh-free convolutions ensure consistent, mesh-independent convergence, also in the presence of noisy inputs. Moreover, our empirical findings suggest that our multiscale training approach can also be coupled with standard convolutions – positioning our multiscale training approach as an alternative to standard SGD. This approach is well-suited for high-resolution and 3D learning tasks where scalability and precision are crucial. Although the current implementation of FFT-based convolutions like our MFCs is slower than standard convolutions, their theoretical FLOPs count is competitive. We hope that the theoretical merits discussed in this paper will inspire the development of efficient implementations of such methods.
Appendix A Background on the solution of Parabolic Equations
@@ B.1 Proof to Lemma 1
To prove Lemma 1, we further analyze Example 1 by asking: how does the gradient of the function behave as we downsample the mesh? To this end, we study the linear model in Example 1 for a single input and output channels, with a 1D convolution kernel of size 3. The loss to be optimized is defined as:
@@ B.3 Equivalence between regular convolutions and discretized MFCs on a fixed mesh
The implications of the lemma is that given a standard network it is always possible to replace the convolutions with MFC and remain within the same accuracy, as our experiments show.
@@
The proof for this lemma is straight forward. Notice that the convolution is made from two parts. The first is a spatial component that involves the Fourier transform and the operation in Fourier space and the second is a $1\times 1$ convolution that combines the spacial channels. We prove the lemma for a single channel. For problems with $c$ channels it is possible to concatenate the same structure $c$ times with different parameters.
To obtain any $k\times k$ convolution we need to be able to express the convolution stencil by a basis of $k^{2}$ convolutions that are linearly independent. To be specific, the standard convolution can be written as
@@
where ${\bf e}_{ij}$ represents a convolution with a stencil of zeros and $1$ in the $i,j$ location and ${\bf K}_{ij}$ is the $ij$ entry of ${\bf K}$. For Mesh Free convolution discretize on a regular mesh and using different values of $\alpha_{xy}$ then the solutions of the parabolic PDE are linearly independent (Evans, 1998). Now use any combination the parameters ${\bm{\gamma}}$ and vary $\alpha_{xy}$ over $m\geq k^{2}$ different values. Choose $\tau$ sufficiently small such that it covers the same apparatus of the standard convolution. We obtain $k^{2}$ different spatial convolutions $\widehat{{\bf e}}_{i}$, each corresponds to a different tilting parameter $\alpha_{xy}$. Since the new set of convolutions are linearly independent they form a basis for any $m$ convolution. In particular, it is possible to minimize the over-determined linear system for the coefficients ${\bf c}$ such that
@@ Appendix C Computation of #WU within a multiscale framework
Definition 1 (Working Unit (WU)).
@@
We now show how to measure the computational complexity in terms of #WU for the three training strategies SGD, Multiscale-SGD, and Full-Multiscale-SGD. For the Multiscale-SGD approach, all computations happen on the finest mesh (size $h$), while for Multiscale-SGD and Full-Multiscale-SGD, the computations are performed at 4 resolutions $(h,2h,4h,8h)$. The computation of running the network on half resolution is 1/4 of the cost, and every coarsening step reduces the work by an additional factor of 4. In the Equations below, we denote use the acronyms MSGD for Multiscale-SGD, and FMSGD for Full-Multiscale-SGD From equation 6, we have,
@@
Alternately, seeing an equivalent amount of data, doing these same computations on the finest mesh with SGD, the #WU per iteration is given by, $N_{0}\times 1+N_{1}\times 1+N_{2}\times 1+N_{3}\times 1$ images in one iteration (where each term is computed at the finest scale). With $N_{1}=2N_{0}$, $N_{2}=4N_{0}$, and $N_{3}=8N_{0}$, the total #WU for $I$ iterations in this case, becomes,
@@
where $I_{h},I_{2h},I_{4h}$ and $I_{8h}$ represent the number of training iterations at each scale. Choosing $I$ iterations at the coarsest scale with $I=I_{8h}=2I_{4h}=4I_{8h}=8I_{h}$ and $N_{1}^{r}=2N_{0}^{r}$, $N_{2}^{r}=4N_{0}^{r}$, and $N_{3}^{r}=8N_{0}^{r}$ for each $r\in\{h,2h,4h,8h\}$, the total #WU for Full-Multiscale-SGD become,
#$\text{{\#WU\ }}_{\text{FMSGD}}=\frac{37}{16\cdot 8}N_{0}^{h}I+\frac{17}{32\cdot 4}N_{0}^{2h}I+\frac{7}{64\cdot 2}N_{0}^{4h}I+\frac{1}{64}N_{0}^{8h}I#$ (47)
Finally, choosing $N_{0}^{2^{j}h}=2^{j}N_{0}$, total #WU simplifies to,
#$\text{{\#WU\ }}_{\text{FMSGD}}=\frac{115}{128}N_{0}I\approx 0.90N_{0}I#$ (48)
Thus, it is roughly $16$ times more effective than using SGD.

### v2

@@ 5 Experimental Results
In this section, we evaluate the empirical performance of our training strategies: Multiscale-SGD (Algorithm 1) and Full-Multiscale-SGD (Algorithm 2), compared to the standard CNN training with SGD. We demonstrate the broad application of our training strategies on architectures like ResNet (He et al., 2016), UNet (Ronneberger et al., 2015), and ESPCN (Shi et al., 2016) on a wide variety of tasks ranging from image denoising, deblurring, inpainting, and super-resolution.
For each of the tasks, we compare a standard network along with their MFC versions, where we replace standard convolutions with our mesh-free convolutions (Section 3.1). We denote these modified architectures as e.g., MFC-ResNet, MFC-UNet, MFC-ESPCN, etc. Additional details on experimental settings, hyperparameters, architectures, and datasets are provided in Appendix E.
@@
Metrics. To address our questions, we focus on performance metrics (e.g., MSE or SSIM) and the computational effort for each method. As a baseline, we train the problem on a single input resolution. To be unbiased to implementation and software limitations, we define a work unit (#WU ) as one application of the model on a single image at the finest mesh (i.e., original resolution). In multiscale training, this unit decreases by a factor of 4 with each downsampling by a factor of 2. We then compare #WU  across training strategies.
Blur Estimation. We begin by solving a quadratic problem to validate the theoretical concepts discussed in Sections 2.2 and 3.2. To this end, we estimate a blurring kernel, ${{\bf G}}$, directly from data (see Nagy & Hansen (2006) for the definition). We assume data ${\bf y}$ of the form ${\bf y}={\bf G}\star{\bf u}+{\bf z}$, where ${\bf u}$ is the original (unblurred) data and ${\bf z}$ is noise. The standard MSE loss is used to estimate ${\bf G}$.
We measure computational work using #WU by running 200 iterations on the fine mesh with SGD, Multiscale-SGD, and Full-Multiscale-SGD, and early-stopping when the validation loss plateaus. Table 3 shows that using Multiscale-SGD yields significant savings, and Full-Multiscale-SGD offers even further savings. We discuss additional details for this experiment in Section D.1.
@@
Image Denoising. Here, one assumes data of the form ${\bf u}^{h}=t{\bf y}^{h}+(1-t){\bf z}$, where ${\bf y}^{h}$ is some image on a fine mesh $h$ and ${\bf z}\sim\mathcal{N}(0,{\bf I})$ is the noise. The noise level $t\in[0,1]$ is chosen randomly. The goal is to recover ${\bf y}^{h}$ from ${\bf u}^{h}$. The loss to be minimized is $loss({\bm{\theta}})={\frac{1}{2}}{\mathbb{E}}_{{\bf y}^{h},{\bf u}^{h},t}\|f({\bf u}^{h},t,{\bm{\theta}})-{\bf y}^{h}\|^{2}$. In Table 4, Multiscale-SGD and Full-Multiscale-SGD significantly improve training compute (as measured by #WU ) while maintaining the quality of the reconstructed denoised image on the STL10 test dataset, measured by MSE. Additional experiments on the CelebA dataset for this task are provided in Section D.2 (see, Table 5).
Table 4: Comparison of different training strategies, SGD, Multiscale-SGD and Full-Multiscale-SGD, over different standard convolution and MFC-based networks across various tasks such as denoising, deblurring, inpainting, and super-resolution.
Image Denoising, MSE ($\downarrow$)
Training strategy #WU ($\downarrow$) UNet ResNet MFC-UNet MFC-ResNet
SGD (Single-scale) 480,000 0.1918 0.1629 0.1800 0.1862
@@
Image Deblurring, MSE ($\downarrow$)
UNet ResNet MFC-UNet MFC-ResNet
SGD (Single-scale) 480,000 0.1510 0.1189 0.1324 0.1063
Multiscale-SGD (Ours) 74,000 0.1458 0.1211 0.1239 0.1057
Full-Multiscale-SGD (Ours) 28,750 0.1561 0.1557 0.1257 0.1191
Image Inpainting, SSIM ($\uparrow$)
UNet ResNet MFC-UNet MFC-ResNet
SGD (Single-scale) 480,000 0.8840 0.8820 0.9161 0.8745
Multiscale-SGD (Ours) 74,000 0.8627 0.8463 0.8828 0.8686
Full-Multiscale-SGD (Ours) 126,000 0.8599 0.8643 0.8749 0.8629
Image Super-resolution, SSIM ($\uparrow$)
@@
SGD (Single-scale) 30,000 0.8491 0.8366 0.8229 0.8451
Multiscale-SGD (Ours) 4,625 0.8138 0.8105 0.7988 0.8300
Full-Multiscale-SGD (Ours) 7,875 0.8042 0.8013 0.7866 0.8173
Image Deblurring. Here, one assumes data of the form ${\bf u}^{h}={\bf K}^{h}{\bf y}^{h}+{\bf z}$, where ${\bf y}^{h}$ is some image on mesh $h$, ${\bf K}^{h}$ is the blurring kernel and ${\bf z}$ is the noise. The goal is to recover ${\bf y}^{h}$ from blurred ${\bf u}^{h}$. The loss to be minimized is $loss({\bm{\theta}})={\frac{1}{2}}{\mathbb{E}}_{{\bf y}^{h},{\bf u}^{h}}\|f({\bf u}^{h},{\bm{\theta}})-{\bf y}^{h}\|^{2}$. In Table 4, our Full-Multiscale-SGD significantly accelerates training while maintaining the quality of the reconstructed deblurred image on the STL10 dataset, measured by MSE.
Image Inpainting. Here, one assumes data of the form ${\bf u}^{h}={\bf M}^{h}{\bf y}^{h}+{\bf z}$, where ${\bf y}^{h}$ is a complete image on mesh $h$, ${\bf M}^{h}$ is the image corruption operation and ${\bf z}$ is the noise. The goal is to recover ${\bf y}^{h}$ from incomplete noised data ${\bf u}^{h}$. The loss to be minimized is $loss({\bm{\theta}})={\frac{1}{2}}{\mathbb{E}}_{{\bf y}^{h},{\bf u}^{h}}\|f({\bf u}^{h},{\bm{\theta}})-{\bf y}^{h}\|^{2}$. In Table 4, our Full-Multiscale-SGD significantly accelerates training while maintaining the quality of the reconstructed inpainted image on the CelebA dataset, measured by SSIM.
Image Super-resolution. Here, we aim to predict a high-resolution image ${\bf u}^{h}$ from a low-resolution image ${\bf y}^{l}$, which is a downsampled version of ${\bf u}^{h}$. The downsampling process is modeled as ${\bf y}^{l}={\bf D}{\bf u}^{h}+{\bf z}$, where ${\bf D}(\cdot)$ is a downsampling operator (e.g., bicubic), and ${\bf z}$ is noise. The goal is to reconstruct ${\bf u}^{h}$ using a neural network $f({\bf y}^{l},{\bm{\theta}})$. The loss function is $\mathcal{L}({\bm{\theta}})=-\mathbb{E}_{{\bf y}^{l},{\bf u}^{h}}\left[\text{SSIM}\left(f({\bf y}^{l},{\bm{\theta}}),{\bf u}^{h}\right)\right].$ In Table 4, our Full-Multiscale-SGD significantly accelerates training while maintaining image quality on the Urban100 dataset, measured by SSIM.
Detailed experimental settings and visualizations for each of the above tasks have been provided in Appendix E and Appendix F, respectively. Furthermore, we provide additional experiments with deeper networks and other type of convolutions in Sections D.3 and D.4, respectively. We also compare runtimes between standard convolutions and MFCs in milliseconds and FLOPs across different image and kernel sizes and runtime per training iteration for SGD and Multiscale-SGD strategies in Section D.5.
6 Conclusion
In this paper, we introduced a novel approach to multiscale training for deep neural networks, addressing the limitations of standard mesh-based convolutions. We showed that theoretically, standard convolutions may not converge as $h\rightarrow 0$, hindering performance. To solve this, we proposed mesh-free convolutions inspired by parabolic PDEs with trainable parameters, ensuring consistent, mesh-independent convergence. Our multiscale training approach can also be coupled with standard convolutions, positioning it as an alternative to standard SGD. This approach is well-suited for high-resolution and 3D learning tasks where scalability and precision are crucial. While FFT-based MFCs are currently slower than standard convolutions, their theoretical FLOPs count is competitive, motivating future efficient implementations.
Impact Statement
This paper presents a method for improving the scalability and efficiency of CNN training through multiscale optimization. Our approach reduces training compute costs while maintaining the performance of deep networks across various tasks, which is particularly useful for high-resolution learning tasks. Additionally, our findings suggest that multiscale training can be integrated with standard convolutions across various network architectures as an alternative to conventional SGD. This paper presents work with a goal to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Background on the solution of parabolic partial differential equation
@@ Appendix B Proofs
To prove Lemma 2.1, we further analyze Example 1 by asking: how does the gradient of the function behave as we downsample the mesh? To this end, we study the linear model in Example 1 for a single input and output channels, with a 1D convolution kernel of size 3. The loss to be optimized is defined as:
@@ B.3 Equivalence between regular convolutions and discretized MFCs on a fixed mesh
The implications of the lemma is that given a standard network, it is always possible to replace the convolutions with MFC and remain within the same accuracy, as our experiments show.
@@
The proof for this lemma is straightforward. Notice that the convolution is made of two parts. The first is a spatial component that involves the Fourier transform and the operation in Fourier space, and the second is a $1\times 1$ convolution that combines the spatial channels. We prove the lemma for a single channel. For problems with $c$ channels, it is possible to concatenate the same structure $c$ times with different parameters.
To obtain any $k\times k$ convolution, we need to be able to express the convolution stencil by a basis of $k^{2}$ convolutions that are linearly independent. To be specific, the standard convolution can be written as
@@
where ${\bf e}_{ij}$ represents a convolution with a stencil of zeros and $1$ in the $i,j$ location and ${\bf K}_{ij}$ is the $ij$ entry of ${\bf K}$. For Mesh Free convolution, discretize on a regular mesh and using different values of $\alpha_{xy}$, then the solutions of the parabolic PDE are linearly independent (Evans, 1998). Now use any combination of the parameters ${\bm{\gamma}}$ and vary $\alpha_{xy}$ over $m\geq k^{2}$ different values. Choose $\tau$ sufficiently small such that it covers the same apparatus of the standard convolution. We obtain $k^{2}$ different spatial convolutions $\widehat{{\bf e}}_{i}$, each corresponds to a different tilting parameter $\alpha_{xy}$. Since the new set of convolutions are linearly independent, they form a basis for any $m$ convolution. In particular, it is possible to minimize the over-determined linear system for the coefficients ${\bf c}$ such that
@@ Appendix C Computation of #WU within a multiscale framework
Definition C.1 (Working Unit (WU)).
@@
We now show how to measure the computational complexity in terms of #WU for the three training strategies SGD, Multiscale-SGD, and Full-Multiscale-SGD. For the Multiscale-SGD approach, all computations happen on the finest mesh (size $h$), while for Multiscale-SGD and Full-Multiscale-SGD, the computations are performed at 4 resolutions $(h,2h,4h,8h)$. The computation of running the network on half resolution is 1/4 of the cost, and every coarsening step reduces the work by an additional factor of 4. In the Equations below, we denote use the acronyms MSGD for Multiscale-SGD, and FMSGD for Full-Multiscale-SGD From equation 2.1, we have,
@@
Alternatively, seeing an equivalent amount of data, doing these same computations on the finest mesh with SGD, the #WU per iteration is given by, $N_{0}\times 1+N_{1}\times 1+N_{2}\times 1+N_{3}\times 1$ images in one iteration (where each term is computed at the finest scale). With $N_{1}=2N_{0}$, $N_{2}=4N_{0}$, and $N_{3}=8N_{0}$, the total #WU for $I$ iterations in this case, becomes,
@@
where $I_{h},I_{2h},I_{4h}$ and $I_{8h}$ represent the number of training iterations at each scale and $N_{1}^{r}=2N_{0}^{r}$, $N_{2}^{r}=4N_{0}^{r}$, and $N_{3}^{r}=8N_{0}^{r}$ for each $r\in\{h,2h,4h,8h\}$ with $N_{0}^{2^{j}h}=2^{j}N_{0}$. For the denoising and deblurring tasks, we chose $I$ iterations at the coarsest scale with $I=I_{8h}=2I_{4h}=4I_{2h}=8I_{h}$. The total #WU for Full-Multiscale-SGD for these two tasks is given by,
#$\text{{\#WU\ }}_{\text{FMSGD}}=\frac{37}{16\cdot 8}N_{0}I+\frac{17}{32\cdot 4}2N_{0}I+\frac{7}{64\cdot 2}4N_{0}I+\frac{1}{64}8N_{0}I=\frac{115}{128}N_{0}I\approx 0.90N_{0}I#$ (45)
Thus, it is roughly $16$ times more effective than using SGD. For both denoising and deblurring, we chose $N_{0}=16$ and $I=2000$.
Similarly, for the inpainting and super-resolution tasks, we chose $I$ iterations at the coarsest scale with $I=I_{8h}=I_{4h}=I_{2h}=I_{h}$. We observed that a larger number of iterations per level were required for these tasks to achieve a similar accuracy as SGD. The #WU for Full-Multiscale-SGD for these three tasks is given by,
#$\text{{\#WU\ }}_{\text{FMSGD}}=\frac{37}{16}N_{0}I+\frac{17}{16}N_{0}I+\frac{7}{16}N_{0}I+\frac{1}{8}N_{0}I=\frac{63}{16}N_{0}I\approx 3.94N_{0}I#$ (46)
Thus, it is roughly $3.8$ times more effective than using SGD. For inpainting and super-resolution, we chose $N_{0}=16$ and $I=2000$, and $N_{0}=8$ and $I=250$, respectively.